OpenStudy (anonymous):

If a right triangle has a hypotenuse of 12 cm and one leg that measures 9 cm, find the length of the other leg.

OpenStudy (anonymous):

\[9^2 + x^2 = 12^2\]

OpenStudy (anonymous):

sqrt(12^2 - 9^2) = sqrt (144-81) = sqrt 63

OpenStudy (anonymous):

By first using Pythagora.

OpenStudy (anonymous):

= 7.94

OpenStudy (anonymous):

thank you